# Self Supervised Networks for Learning Latent Space Representations of Human Body Scans and Motions

### v1

@@ 1 Introduction
1.1 Motivation
The examination of human body shapes holds significant importance in various fields such as computer vision, graphics, and virtual reality. The key deliverables in this areas include:
@@ 1.1 Motivation
    Fast algorithms for clustering and classification of (large scale) data of unregistered human body scans;
  2. 2.
    Efficient computational frameworks for accurate motion transfer, motion interpolation and motion extrapolation;
  3. 3.
    Generative modelling via empirical distributions learned from (large scale) unregistered human body scan data bases.
Numerous studies [2, 16, 29, 19, 24] have focused on addressing these challenging computational tasks by parameterizing the space of human body shapes using low-dimensional latent space models. However, several difficulties persist with these latent space parameterizations.
One key challenge consists in the non-uniqueness of the raw shape data, i.e., the same human body shape can be represented by infinitely many different mesh representation. In the continous (smooth) setting this non-uniqueness is encoded by defining two shapes to be the same if they only differ by a reparametrization. The analogue in the discrete setting is to consider two human body scans (represented as simplicial complexes) to be the same if they only differ by a remeshing operation. This non-uniqueness makes the task of retrieving latent code representations from unregistered data significantly more challenging. In the quest for faithful latent space representation of the space of human body shapes it is thus paramount to develop mesh invariant latent space representations from raw body scans with minimal mesh preprocessing and computational demands.
A second key challenge consists of the fact that the linear geometry of a given latent space representation will not accurately describe the non-linear nature of human body movements, which brings up the challenging task of accurately capturing the geometry of human motions within latent spaces and thereby obtaining a faithful computational framework based on these latent space representations.
1.2 Contributions
Towards these ends, we propose two self-supervised deep learning models which can be used in conjunction to address these demands:
  1. 1.
    First, we introduce VariShaPE (Varifold Shape Parameter Estimator), a self-supervised neural network model for real-time mesh invariant latent space encoding of human body scans;
@@ 1.2 Contributions
We note that, as demonstrated in Section 4, these two models can be used in conjunction to obtain efficient (real time) and accurate algorithms for operations such as motion transfer, generative modelling and 4D-interpolation. Next we describe the benefits of the two methods, as compared to state-of-the-art algorithms, in more details.
Benefits of VariShaPE and Comparison to State-of-the-art Methods:
Finding low-dimensional parametric representation for the statistical analysis of human shapes is an active research area [2, 16, 29], that dates back to the groundbreaking work on 3D Morphable Models for faces by Blanz et al. [4]. In the particular case of modelling for whole body scans, common approaches include skeleton based models in combination with physically motivated transformation models. Among those, one of the most widely used approaches, is the SMPL latent space [19], which is learned from thousands of 3D body scans; see also [26, 24] for recent extensions that further enhance this model by adding eg. articulated hands and face expressions.
Our first contribution, VariShaPE, deals with the challenging task of estimating such latent space representations given raw data, i.e., to retrieve the best latent parameter to fit a given unregistered scan of a human body. Standard methods for this retrieval task are usually optimization based and often require manual intervention in order to lead to satisfactory results [19]; this includes spectral based methods such as the functional map framework [25], minimizing the Hausdorff or Chamfer distance [9, 13], or methods based on Riemannian geometry and elastic shape analysis [15]. In contrast to these methods our model employs a self-supervised network approach making use of concepts of geometric measure theory [18], which allows us to incorporate shape invariances directly in the network architecture and thereby significantly boost the performance of the corresponding algorithms.
A further key benefit of the proposed VariShaPE model is that it can be trained on (applied to, resp.) raw scans without any additional mesh preprocessing. This is precisely due to the aforementioned geometric measure theory inspired architecture, which is specifically designed to ensure complete mesh invariance, i.e., our model will produce (approximately) the same latent code representations of scans that represent the same human body invariant to the mesh structure or scanning noise which appears in real data. An additional benefit of our framework is the fast training time as compared to other network-based approaches in human shape analysis. This again stems from our use of geometrically motivated constructions in the network, which allows us to encode the invariances of the data directly in the corresponding architecture. Thereby we are able to significantly reduce the amount of trainable parameters and thus require an order of magnitudes lower amount of training time as compared to previous approaches such as LIMP [8], ARAPReg [17], or 3D-Coded [13]. Finally we present results that show our latent code retrieval framework is competitive with the state of the art 3D-Coded [13] method for shape registration while requiring orders of magnitude less computational time, c.f Table 1 in Section 3.1. Additionally we provide experiments to highlight our models superior robustness to mesh resampling, c.f Table 2 in Section 3.1.
Benefits of MoGeN and Comparison to State-of-the-art Methods:
Once data (eg. human body scans) is embedded in a linear (Euclidean) latent space, one directly obtains a corresponding framework for statistical analysis: operations such as interpolation and extrapolation can be simply defined by performing these operations in the linear latent space. In the context of human body scans, this raises the issue of whether the linear geometry of the latent space accurately describes the complex movements and deformations appearing in such data, see also [21, 20, 30] for a discussion on this matter in the context of latent space interpolation for general machine learning applications.
Indeed, the performance of such a naive approach is rather limited in the context of large movements of human bodies. To address these issues multiple (physically motivated) deformation energy losses have been introduced in the training phase, thereby altering the corresponding latent space: this includes geodesic distances [8], the As-Rigid-As-Possible (ARAP) energy [17, 22], or volumetric constraints [3]. See also [31, 10] for an approach using manifold regularization of learned pose spaces. These geometric quantities, however, significantly increase the total training costs of those approaches. More importantly, they are still based on user-defined assumption to model the movement of human shapes. Our work differs in two different aspects from these approaches: first we keep the latent space unchanged, but instead equip it with a different non-linear geometry. Secondly, we do not make any assumptions on the physics behind the deformations of human body motions, but instead learn them in a purely data-driven approach using 4D training data. In our experiments we argue that this data driven approach leads to a significantly improved performance in terms of interpolation and extrapolation accuracy. A further key benefit of this methodology is that it allows us to capture more complex motions of human bodies, including those which reach the ”boundary” of feasible human motions.
Future work and Choice of Latent Space Model:
In summary, we have proposed in this paper new self-supervised models for learning not only latent space representations of human body shape and pose but also the underlying geometry of the latent space itself. The latter one is achieved through the introduction of an additional lifted space in which observed body pose motions get unfolded to straight paths. We have further shown how these two frameworks can be used to perform common tasks such as motion interpolation, extrapolation and transfer, and to estimate generative models in the space of human body pose and shape.
@@
In our experimental section, we demonstrate the capabilities of our models using data from the DFAUST dataset and by employing the SMPL model (A Skinned Multi-Person Linear body model) [19] as the corresponding latent space representation. We want to emphasize, that our framework is not tied in any way to this particular choice, but could be directly applied to any other existing latent space representation for human bodies, such as SMPL-X [26], STAR [24], BLISS [23] or the elastic shape analysis basis representation used in [15, 28]. While we aim to extend our framework in future work to these latent spaces, we decided to focus in the experiments of this present paper on the SMPL representation, due to its convenient implementation in PyTorch and its wide use in the human body shape community. In one of our ablation studies, cf. Appendix A, we demonstrate a similar performance using the elastic shape analysis basis setting as an alternative latent space representation; in future work we plan to explore the use of some of the recent SMPL extensions, such as the above mentioned STAR [24].
@@ 2 Methods Section
In this section we will introduce the main methodological contribution of the present work. We start by introducing the space of triangular meshes $\mathcal{S}$.
A mesh with $m_{v}$ vertices and $m_{f}$ faces is described by an ordered set of vertices $\mathcal{V}=\{v_{i}\}^{m_{v}}_{i=1}\subseteq\mathbb{R}^{3}$ and an ordered set of faces described by the indices of the vertices of the face =$\mathcal{F}=\{f_{j}\}^{m_{f}}_{j=1}\subseteq(1,2,...,m_{v})\times(1,2,...,m_{v})\times(1,2,...,m_{v})$. The space of meshes is then defined as
$\mathcal{S}:=\left\{\left(\{v_{i}\}^{m_{v}}_{i=1},\{f_{j}\}^{m_{f}}_{j=1}\right)\text{ s.t. }m_{v}\in\mathbb{N}\text{ and }m_{f}\in\mathbb{N}\right\}.$
Note, that the elements of $\mathcal{S}$ can differ both in the mesh face connectivity or even in the total number of elements (edges and faces, resp.).
Remark 1 (Non-uniqueness of mesh representations).
It is important to note, that the representation of a human body as an element of $\mathcal{S}$, i.e., as a triangulated mesh, comes with several ambiguities: many meshes can represent the same human scan and thus one has to compare these representations up to shape preserving transformations. These shape preserving transformations are given by translation, rotation and (most importantly) general remeshing operations such as subdivison algorithms.
In the exploration of human body surfaces, a prevalent strategy for tackling the dimension reduction problem involves adopting a latent space model. Such a model is essentially a mapping $F$ from a latent space $\mathcal{L}^{F}\subset\mathbb{R}^{n}$ to the space of triangulated surfaces $\mathcal{S}$. Various construction methods, such as data-driven basis construction or skinned skeletal models, can be employed to build such latent spaces, see [19, 26, 24, 23, 28] and the references therein. The first focus of this paper is to solve the inverse problem associated with a fixed latent space model, i.e, for an arbitrary mesh $q$ and a given latent space model $F:\mathcal{L}^{F}\to\mathcal{S}$ we aim to find $v\in\mathcal{L}^{F}$ such that $F(v)$ and $q$ represent the same shape. We will refer to this as the latent code retrieval problem.
To deal with the aforementioned non-uniqueness in the mesh representation of a human body shape, we will start by introducing a concept from geometric measure theory, which will allow us to define a fixed-dimension, mesh invariant feature vector representation, which we can then used as the input to train a fully connected neural network, $\Psi_{\theta}$, parameterized by weights $\theta=\{\theta_{i}\}$, mapping this feature vector representation into the latent space $\mathcal{L}^{F}$.
2.1 The Varifold Gradient as a Shape Descriptor
Many works in the field of human body analysis, including the aforementioned references [19, 9, 13], have relied on the well-known Chamfer or variants of the Hausdorff distance to define loss functions for surface reconstruction or estimation of latent representations. However, these typically result in loss functions that are not differentiable at every point. Alternatives, which arguably lead to better regularity and stability properties, can be obtained via representations of surfaces in measure spaces equipped with kernel metrics [7] (also known as maximum mean discrepancy in statistics [11]), among which the varifold framework has recently seen several applications to the field [1, 27, 15]. In the present work, we will leverage varifolds for two specific purposes:
@@ 2.1 The Varifold Gradient as a Shape Descriptor
    to obtain a reliable notion of surface distance to evaluate and benchmark our proposed algorithms and, more importantly,
@@
    to obtain a mesh invariant feature representation for the proposed VariShaPE method for solving the latent code retrieval problem.
Let us briefly recap the primary features of the varifold framework for surface meshes, referring to [18, 7] for more detailed presentations:
Definition 1 (Varifold distance).
Given two triangulated surfaces $q,q^{\prime}\in\mathcal{S}$ denote by $\{f_{1},\ldots,f_{m}\}$ and $\{f^{\prime}_{1},\ldots,f_{m^{\prime}}\}$ the respective set of non-degenerate triangular faces in $q$ and $q^{\prime}$. The (squared) varifold distance is then computed by examining all possible pairs of faces in the following way:
$d_{\operatorname{Var}}(q,q^{\prime})^{2}=$ $\sum_{i=1}^{m}\sum_{j=1}^{m}\rho(|c_{f_{i}}-c_{f_{j}}|)\frac{(n_{f_{i}}\cdot n_{f_{j}})^{2}}{|n_{f_{i}}||n_{f_{j}}|}+\sum_{i=1}^{m^{\prime}}\sum_{j=1}^{m^{\prime}}\rho(|c_{f_{i}}-c_{f^{\prime}_{j}}|)\frac{(n_{f^{\prime}_{i}}\cdot n_{f^{\prime}_{j}})^{2}}{|n_{f^{\prime}_{i}}||n_{f^{\prime}_{j}}|}$
$-2\sum_{i=1}^{m}\sum_{j=1}^{m^{\prime}}\rho(|c_{f_{i}}-c_{f^{\prime}_{j}}|)\frac{(n_{f_{i}}\cdot n_{f^{\prime}_{j}})^{2}}{|n_{f_{i}}||n_{f^{\prime}_{j}}|}$ (1)
where for a face $f$ (of $q$ or $q^{\prime}$), $c_{f}$ denotes its barycenter, $n_{f}$ is the normal vector of length equal to the area of $f$ and where $\rho$ is a given kernel function. For notational purposes we denote the squared distance from a fixed shape $q$ by $D_{q}(q^{\prime}):=d_{\operatorname{Var}}(q,q^{\prime})^{2}$.
Note, that the above definitons does not assume that the surfaces $q$ and $q^{\prime}$ are registered, i.e., their number of vertices and triangles may differ and no prior correspondences are known.
Remark 2 (Choice of kernel function).
In this work we choose the kernel function $\rho$ to be a centered Gaussian which variance can be interpreted as the spatial scale of the metric. In particular, we choose $\rho(|c-c^{\prime}|)=e^{-\frac{|c-c^{\prime}|^{2}}{\sigma^{2}}}$ with $\sigma=0.025$. Note that many other choices of kernels on the face centers and normals are also possible, c.f. [18].
@@
Furthermore, we point out that the squared distance in (1) is differentiable with respect to the vertex positions of $q$ and $q^{\prime}$ provided the kernel function $\rho$ is smooth enough. The gradients can be easily computed either explicitly from (1) or via automatic differentiation, with high numerical efficiency on a GPU due to the very parallelizable nature of this computation. This allows us to introduce the gradient of the varifold distance, which we will later use as a fixed dimension shape descriptor:
Definition 2 (VariGrad Shape Representation).
Given a fixed template mesh $\bar{q}\in\mathcal{S}$ with $M$ vertices we define
$\operatorname{VariGrad}:\begin{cases}\mathcal{S}&\to\mathbb{R}^{3M}\text{ given by }\\ q&\mapsto\nabla D_{q}(\bar{q}+(\cdot))(0).\end{cases}$
For our purposes, the main advantage of the VariGrad representation is given by the following two properties:
  1. 1.
    $\operatorname{VariGrad}(q)$ is a vector field on $\bar{q}$ and thus has fixed dimension $3M$ independently of the dimension of the input mesh $q$. Intuitively, it encodes the positioning of $q$ with respect to $\bar{q}$; a formal analogy can be drawn with the gradient of a squared Riemannian distance that corresponds to the logarithm map of the manifold.
  2. 2.
    Due to the properties of varifold metrics it is essentially invariant to resampling or mesh changes in the surface represented by $q$ as well as certain amount of imaging noise, which is showcased in the example of Figure 1.
Thus it can be leveraged as an effective feature representation for latent space neural network learning models, as has been recently shown in the context of curves and shape graphs, see [14].
@@
We are now ready to introduce the proposed model for solving the latent code retrieval problem. Inspired by the recent work on curves and shape graphs [14] the underlying idea is to combine the mesh invariant VariGrad representation with a fully connected neural network, $\Psi_{\theta}$, which is trained to learn the mapping from this fixed dimension feature vector space into the given latent space model $F:\mathbb{R}^{n}\supset\mathcal{L}^{F}\to\mathcal{S}$.
Figure 2: Illustration of the proposed VariShaPE framework: The framework utilizes two non-trainable latent space models: The first, $G$, is utilized to construct feature vectors from unregistered meshes using the VariGrad operator defined above. The second, $F$, parameterizes the space of human bodies. Given an input mesh $q$ with arbitrary mesh structure, we compute the VariGrad shape descriptor with respect to the parameters of the model $G$, $\operatorname{VariGrad}_{G}(q)$. These parameterization invariant features are passed through a trainable MLP with parameters $\theta$ to predict a latent space representantion of $q$ with respect to the target latent space model $F$. As one can see in the example presented in this figure, our framework is robust to imaging noise, i.e., holes and small missing parts stemming eg. from imaging noise are not significantly influencing the reconstruction procedure.
The drawback of this approach in the context of human body surfaces is that the template requires a large number of vertices to be sufficiently expressive, thereby producing a high dimensional feature vector representation and thus a higher computational complexity for training. To circumvent this difficulty we will construct a feature space utilizing a second reduced dimension latent space model $G:\mathbb{R}^{m}\supset{\mathcal{L}}^{G}\to\mathcal{S}$ and define a shape descriptor based on this model as follows:
Definition 3 (Latent space constrained VariGrad Shape Representation).
For a given latent space model $G:\mathbb{R}^{m}\supset\mathcal{L}^{G}\to\mathcal{S}$ we define the VariGrad shape descriptor as
$\operatorname{VariGrad}_{G}:\begin{cases}\mathcal{S}&\to\mathcal{L}^{G}\subset\mathbb{R}^{m}\text{ given by }\\ q&\mapsto\nabla(D_{q}\circ G)(\overrightarrow{0}).\end{cases}$
Note that the unrestricted Varigrad of Definition 2 corresponds to the particular case in which $G(u)=\bar{q}+u$. The advantage of introducing a more general latent space model $G$ lies in the fact that we only calculate the gradient w.r.t. the variables of a lower-dimensional latent space. Thus the latent space constrained VariGrad operator produces shape descriptors with a potentially (order of magnitudes) lower dimension than that of Definition 2 while still remaining insensitive to mesh resampling. It is important to note, that this additional latent model $G$ can be chosen independently of the targeted latent space representation $F$; it is merely a computational tool and (beyond computational accuracy) it is not influencing the final latent code retrieval. This is supported by our experiments, see Section 3 and the ablation study in the supplementary material, in which we choose different latent models for $F$ and $G$ and justify our particular choices.
{mdframed}
The latent code retrieval algorithm, which we propose can then be summarized as follows
  • for an arbitrary mesh representing a human body, $q$, we consider the latent space constrained VariGrad shape descriptor (see Definition 3) for a frozen latent space model $G:\mathbb{R}^{m}\to\mathcal{S}$ to define an $m$-dimensional mesh invariant descriptor of $q$.
  • These representations are then used as inputs to an MLP $\Psi_{\theta}$, which is trained to learn the desired latent code representation $v\in\mathcal{L}^{F}$.
  • We train the model so that $F(v)$ and $q$ represent the same shape, i.e., we consider a Loss function of the form
    $\operatorname{Loss}(F(v))=D_{q}(F(v)),$ (2)
    where $\operatorname{dist}$ is an appropriate similarity measure on the space of triangulated surfaces.
A schematic of this framework is presented in Figure 2. It remains to specify the loss function $\operatorname{Loss}$, for which there are several natural choices: if the available training data has the same mesh structure as the outputs of the model $F$ (with consistent point correspondences) we can train the model to simply minimize the mean vertex error (MVE) between $F(v)$ and $q$, i.e.,
$\operatorname{Loss}^{\operatorname{MVE}}(F(v)):=\frac{1}{k}\sum_{i=1}^{k}\|F(v)_{i}-q_{i}\|^{2}.$ (3)
Alternatively if such training data is not available, then the Chamfer distance and Varifold distance (as defined in (1)) both offer differentiable loss functions for training data without point to point correspondences to the outputs of $F$. For the experiments presented in this paper, we train with a subset of the DFAUST dataset (c.f Section 3) which falls under the former category, and thus we present results of the model trained with the mean vertex error as the loss function of the model. An ablation study for these other choices of loss functions is given in Section B.
Remark 3.
@@
In recent years, several other frameworks for sampling invariant feature vector representations of geometric data have been proposed. In the context of point clouds, PointNet [6] and Dynamic Graph Convolutional Neural Networks (DGCNN) [32], propose frameworks for constructing global features with fixed dimension representing this data. As we will see in the experiments section, our approach leads to a significantly improved performance for dealing with the resampling (remeshing, resp.) problem as compared to the PointNet and DGCNN architectures.

### v2

@@ Abstract
keywords:
human body shapes, latent representation learning, varifold, motion interpolation and transfer, generative models
@@ 1 Introduction
The examination of human body shapes holds significant importance in various fields such as computer vision, graphics, and virtual reality.This paper introduces new self-supervised deep learning models which, together, can be used to tackle some of the most common challenges in the field of human body shape analysis under minimal requirements on mesh preprocessing. Our main contributions are two-fold:
@@
    First, we introduce VariShaPE (Varifold Shape Parameter Estimator), a self-supervised neural network model for real-time latent space encoding of unregistered human body scans;
@@
Direct applications of the proposed algorithms include:
  1. 1.
    A learning based framework for accurate motion transfer, motion interpolation and motion extrapolation;
  2. 2.
    Fast algorithms for clustering and classification of (large scale) datasets of unregistered human body scans;
  3. 3.
    Generative modeling via empirical distributions learned from (large scale) unregistered human body scan data bases.
@@
In our experimental section we demonstrate the capabilities of our framework on the DFAUST dataset and by relying on the the SMPL model (A Skinned Multi-Person Linear body representation) [1] as the corresponding latent space representation. We want to emphasize, that our framework is not tied in any way to this particular latent space model, but could be directly applied to any other existing latent space representation for human bodies, such as SMPL-X [2], STAR [3], BLISS [4] or the elastic shape analysis basis representation used in [5, 6, 7]. We decided to focus the experiments on the SMPL representation, due to its convenient handling in PyTorch and its wide use in the the field of 3D human body shape community. In one of our ablation studies, cf. Section 4.4, we demonstrate a similar performance using the elastic shape analysis basis representation as an alternative latent space; in future work we plan to explore the use of some of the recent SMPL extensions, such as the above mentioned STAR [3].
An additional benefit of our framework is the fast training time as compared to other network-based approaches in human shape analysis. This stems from our use of geometrically motivated constructions in the network, which allows us to encode fundamental invariances directly in the proposed architecture. Thereby we are able to significantly reduce the number of trainable parameters and thus require an order of magnitudes lower amount of training time as compared to previous approaches such as LIMP [8], ARAPReg [9], or 3D-Coded [10].
Acknowledgements and Conflicts of Interest
M.B was partially supported by NSF grant CISE 2426549. N.C and E.H were partially supported by NSF grants NSF CISE-2426550 and NSF DMS-2438562.
The authors certify that they have no conflicts of interest.
2 Related Work
2.1 Latent Space Models for Human Body Scan Representations
Finding low-dimensional parametric representation for the statistical analysis of human shapes is an active research area [11, 12, 13], that dates back to the groundbreaking work on 3D Morphable Models for faces by Blanz et al. [14]. In particular applications to whole human body scans, common approaches include skeleton based apporaches in combination with physically motivated transformation models. Among those, one of the most widely used is the SMPL latent space [1], which is learned from thousands of 3D body scans; see also [2, 3] for recent extensions that further enhance this model by adding e.g. articulated hands and face expressions.
Our first contribution deals with the challenging task of extracting such latent space representations given raw data, i.e., to retrieve the best set of latent parameters to fit a given unregistered scan of a human body. Standard methods for this retrieval task are usually optimization based and often require manual intervention in order to lead to satisfactory results [1]; this includes spectral based methods such as the functional map framework [15], minimizing the Hausdorff or Chamfer distance [16, 10], or methods based on Riemannian geometry and elastic shape analysis [5, 6, 17, 18, 19]. In contrast to these methods, our model employs a self-supervised network approach making use of tools from geometric measure theory [20], which allows us to incorporate shape invariances directly in the network architecture and thereby significantly boost the performance of the corresponding algorithms.
2.2 Interpolation in Latent Spaces
Once data (e.g. human body scans) is embedded into a linear (Euclidean) latent space, one directly obtains a corresponding framework for statistical analysis: operations such as interpolation and extrapolation can be simply defined by performing these operations in the linear latent space. In the context of human body scans, this raises the issue of whether the linear geometry of the latent space accurately describes the complex movements and deformations that can be found, see also [21, 22, 23] for a discussion on this matter in the context of latent space interpolation for general machine learning applications.
Indeed, the performance of such a naive approach is rather limited when it comes to large movements of human bodies. To address these issues, multiple (physically motivated) deformation energy losses have been introduced in the training phase, thereby altering the corresponding latent space: this includes geodesic distances [8], the As-Rigid-As-Possible (ARAP) energy [9, 24], or volumetric constraints [25]. See also [26, 27] for an approach using manifold regularization of learned pose spaces. These geometric constraints, however, significantly increase the total training costs of those approaches. More importantly, they are still based on user-defined assumption to model the movement of human shapes. Our work differs in two different aspects from these approaches: first, we keep the latent space unchanged, but instead equip it with a different non-linear geometry. Secondly, we do not make any assumption on the physics behind the deformations of human body motions, but instead learn them in a purely data-driven approach using 4D training data. In our experiments, we argue that this data driven approach leads to a significantly improved performance in terms of interpolation and extrapolation accuracy.
@@ 3 Methods Section
In the exploration of human body surfaces, a prevalent strategy for tackling the dimension reduction problem involves adopting a latent space model. This model is essentially a mapping $F$ from a latent space $\mathcal{L}\subset\mathbb{R}^{n}$ to the space of parameterized surfaces $\mathcal{S}$ (triangulated meshes, respectively). Various construction methods, such as data-driven basis construction or skinned skeletal models, can be employed to build such latent spaces, see [1, 2, 3, 4, 7] and the references therein. The first focus of this paper is to solve the latent code retrieval problem,, that is, for an arbitrary mesh $q$ and a given latent space model $F:\mathcal{L}\to\mathcal{S}$ we aim to find $v\in\mathbb{R}^{n}$ such that $F(v)$ and $q$ represent the same shape. Whereas, the second contribution aims to employ a data-driven approach to learn the geometry of a given latent space representation based on real human body motions.
Towards the first goal, the latent code retrieval problem, our approach is to train a fully connected neural network, $\Psi_{\theta}$, parameterized by weights $\theta=\{\theta_{i}\}$ mapping a mesh invariant feature vector representation of the input data to $\mathcal{L}$. This approach relies on an appropriate construction of mesh invariant representation of the input and comes with several notable challenges:
@@
    Shape preserving transformations: the representation of a human body as a triangulated mesh comes with several ambiguities, i.e., many meshes can represent the same human scan and thus one has to compare these representations up to shape preserving transformations. These shape preserving transformations are given by translation, rotation and (most importantly) general remeshing operations such as subdivison or upsampling.
@@
    Varying dimension and noise of real mesh data: furthermore, meshes from real applications usually lack a fixed dimension. However, this is a prerequisite for the direct application of the proposed neural network architecture. In addition, these scans can exhibit a significant amount of imaging noise, such as holes or even larger missing parts. Our feature vector representation should remain robust to theses types of imaging noise.
Before we are able to introduce our methods in more details, we will introduce a concept from geometric measure theory, which will be the basis for our mesh invariant feature vector construction.
3.1 The Varifold Distance and its Gradient
Many works in the field of human body analysis, including the aforementioned references [1, 16, 10], have relied on the well-known Chamfer or variants of the Hausdorff distance to define loss functions for surface reconstruction or estimation of latent representations. However, these typically result in loss functions that are not differentiable at every point. Alternatives, which arguably lead to better regularity and stability properties, can be obtained via representations of surfaces in measure spaces equipped with kernel metrics [28], among which the varifold framework has recently seen several applications to the field [29, 30, 5]. In the present work, we also leverage varifolds for two specific purposes: 1) to obtain a reliable notion of surface distance so as to evaluate and benchmark our proposed algorithms and, more importantly, 2) to obtain a mesh invariant feature representation for the VariShaPE architecture of the next section.
Let us briefly recap the primary features of the varifold framework for 3D surface meshes, referring to [31, 20, 28] for more detailed presentations. Assume that $q$ and $q^{\prime}$ are two triangulated 3D surfaces not necessarily registered, i.e. their number of vertices and triangles may differ and no prior correspondences are known. Denoting $\{f_{1},\ldots,f_{m}\}$ and $\{f^{\prime}_{1},\ldots,f_{m^{\prime}}\}$ the respective set of non-degenerate triangular faces in $S$ and $S^{\prime}$, the (squared) varifold distance is computed by examining all possible pairs of faces in the following way:
$d_{\operatorname{Var}}(q,q^{\prime})^{2}=$ $\sum_{i=1}^{m}\sum_{j=1}^{m}\rho(|c_{f_{i}}-c_{f_{j}}|)\frac{(n_{f_{i}}\cdot n_{f_{j}})^{2}}{|n_{f_{i}}||n_{f_{j}}|}+\sum_{i=1}^{m^{\prime}}\sum_{j=1}^{m^{\prime}}\rho(|c_{f^{\prime}_{i}}-c_{f^{\prime}_{j}}|)\frac{(n_{f^{\prime}_{i}}\cdot n_{f^{\prime}_{j}})^{2}}{|n_{f^{\prime}_{i}}||n_{f^{\prime}_{j}}|}$ (1)
$\qquad-2\sum_{i=1}^{m}\sum_{j=1}^{m^{\prime}}\rho(|c_{f_{i}}-c_{f^{\prime}_{j}}|)\frac{(n_{f_{i}}\cdot n_{f^{\prime}_{j}})^{2}}{|n_{f_{i}}||n_{f^{\prime}_{j}}|}$
where for a face $f$ (of $q$ or $q^{\prime}$), $c_{f}$ denotes its barycenter and $n_{f}$ is the normal vector of length equal to the area of $f$. The function $\rho$ defines a kernel function which we take, in this work, to be a centered Gaussian which variance can be interpreted as the spatial scale of the metric. Note that many other choices of kernels on the face centers and normals are also possible, cf. [20].
@@ 3.1 The Varifold Distance and its Gradient
We point out that the squared distance in (1) is differentiable with respect to the vertex positions of $q$ and $q^{\prime}$ provided the kernel function $\rho$ is smooth enough. The gradients can be easily computed either explicitly from (1) or via automatic differentiation, with high numerical efficiency on a GPU due to the very parallelizable nature of this computation. Given a surface $q$, let us denote by $\nabla_{q^{\prime}}d_{\operatorname{Var}}(q,\cdot)^{2}$ the distance gradient with respect to the second surface $q^{\prime}$. The resulting vector field on $q^{\prime}$, due to the properties of varifold metrics, is essentially invariant to resampling or mesh changes in the surface represented by $q$ as well as certain amount of imaging noise, which is showcased in the example of Fig. 1. Thus it can be leveraged as an effective feature representation for latent space neural network learning models, as we explain in the next section.
@@
Figure 2: Illustration of the proposed VariShaPE framework: Given an input mesh $q$ with arbitrary mesh structure VariShaPE returns the estimated Latent Space Representantion. As can been seen on the input mesh our framework is robust to imaging noise, i.e., holes and small missing parts do not present any difficulties for the reconstruction procedure.
@@
In recent years, several frameworks for sampling invariant feature vector representations of geometric data have been proposed. In the context of point clouds, PointNet [32] and Dynamic Graph Convolutional Neural Networks (DGCNN) [33], propose frameworks for constructing global features with fixed dimension representing this data. In the context of curves and shape graphs, the recent VariGrad model of [34] utilizes a data driven method for feature vectors, which are obtained by fixing a template object and computing the gradient of the varifold distance with respect to the vertices of the template object following the setting described in the previous section. As compared to the PointNet and DGCNN architectures this approach led to a significantly improved performance for dealing with the resampling (remeshing, resp.) problem, which is the reason that we chose to adopt it in the present article.
The drawback of this approach in the context of human body surfaces is that the template requires a large number of vertices to be sufficiently expressive. Consequently this produces a high dimensional feature vector representation and thus a higher computational complexity for training. To circumvent this difficulty we employ a second reduced dimension latent space model $G:\mathbb{R}^{m}\to\mathcal{S}$ and instead compute the gradient with respect to the latent codes of $G$. It is important to note this additional latent model $G$ can be chosen independently of the targeted latent space representation $F$; it is merely a computational tool and (beyond computational accuracy) it is not influencing the final latent code retrieval. Indeed in our experiments, see Section 4, we choose different latent model for $F$ and $G$, cf. the ablation study of Section 4.4, where we justify our particular choices.
The latent code retrieval algorithm we propose is thus the following: for an arbitrary mesh representing a human body, $q$, we take the varifold gradient at the template (w.r.t. to the latent codes of $G$) $\nabla_{\vec{0}}d_{\operatorname{Var}}(q,G(\cdot))^{2}$; this yield a feature vector representation in $\mathbb{R}^{m}$, where $m$ is the dimension of the latent model $G$. Note that this representation inherits the low dimensionality from the model $G$ and the invariance to remeshing and robustness to noise from the varifold distance. These representations are then used as inputs to the fully connected network $\Psi_{\theta}$ that learns the latent code $v\in\mathcal{L}$ such that $F(v)$ and $q$ represent (approximately) the same shape. A schematic of this framework is presented in Figure 2.
Figure 3: Illustration of the proposed MoGeN framework: Given two points in latent space (a point and a deformation direction, resp.), the network $f$ maps them into a high dimensional, lifted latent space, where interpolation (extrapolation, resp.) follow the standard linear paths. Subsequently, they are projected back to the original latent space using the network $\pi$, which allows the user to map them into the space of surfaces using the decoder $F$.